You graduate from IU and get a job with an annual salary of $50,000.
a.   (1 point) Your boss calls you into the office and says she is happy with your performance and is going to give you a $1,000 raise. What percentage raise is this for you?
     b.   (1 points) If the Inflation Rate is 2.0%, what is the actual amount of your raise?
Answer #1

(a) Annual Salary = $50000

Raise = $1000

Percentage Raise = i = Raise / Salary * 100% = 1000/50000 * 100% = 2%

(b) According to the Fisher Equation,

(1 + i) = (1 + r) × (1 + π)

where, i = nominal interest rate (rate before inflation) = 2%

r = real interest rate

π = inflation rate = 2%

=> (1+0.02) = (1+r)(1+0.02)

=> r = 0%

Hence, actual amount of raise = 0%

Hence, nominal interest rate = 9.2%